OpenStudy (anonymous):

the catenary has equation y=(e^x+e^-x)/2. graph it and y=x^2 on the same set of coordinate axes. for what values of x are the y values within 1 of each other? within 0.5 of each other?

OpenStudy (anonymous):

\[\begin{cases} y_1=\dfrac{e^x+e^{-x}}{2}\\\\ y_2=x^2 \end{cases}\] The \(y\) values will be within 1 of each other if \(|y_1-y_2|\le1\): http://www.wolframalpha.com/input/?i=%7CCosh%5Bx%5D-x%5E2%7C%3C%3D1 and will be within 0.5 of each other if \(|y_1-y_2|\le0.5\): http://www.wolframalpha.com/input/?i=%7CCosh%5Bx%5D-x%5E2%7C%3C%3D1%2F2